# MCAQ-YOLO: Morphological Complexity-Aware Quantization for Efficient Object Detection with Curriculum Learning

### v1

@@ VII Conclusion
We have presented MCAQ-YOLO, a morphological complexity-aware quantization framework for object detection that combines a hierarchical complexity analyzer, a learned complexity-to-bit mapping network with curriculum learning, and a hardware-aware mixed-precision quantization scheme. Through extensive experiments, we have shown that morphological complexity correlates strongly with quantization sensitivity and that exploiting this relationship yields consistent, statistically significant improvements over uniform and layer-wise quantization baselines.
On a construction safety dataset, MCAQ-YOLO achieves 85.6% mAP@0.5 with an average of 4.2 bits, outperforming uniform 4-bit quantization by 3.5 percentage points and the state-of-the-art HAWQ-V3 by 0.9 percentage points while maintaining a 7.63 compression ratio. These results demonstrate that complexity-aware spatial quantization can deliver substantial accuracy gains over uniform low-bit quantization while preserving efficiency. Cross-dataset evaluations on COCO and Pascal VOC demonstrate 2–3% absolute gains, indicating that the approach generalizes beyond the primary dataset.
Despite its limitations—including additional computational overhead and the absence of a complete theoretical foundation—MCAQ-YOLO illustrates that incorporating domain knowledge about visual complexity into quantization decisions can meaningfully improve the accuracy–compression trade-off. We hope this work serves as a stepping stone toward more general forms of spatially adaptive neural network compression.
Appendix A Supplementary Material
A-A Detailed Hyperparameters
| Parameter | Value |
| Base learning rate | |
| Learning rate schedule | Cosine with warmup |
| Warmup epochs | 3 |
| Weight decay | 0.05 |
| Momentum () | 0.9 |
| Adam | 0.999 |
| Gradient clipping | 1.0 |
| Label smoothing | 0.1 |
| Dropout | 0.0 |
| Batch size | 64 (16 per GPU) |
| Image size | 640640 |
| Multi-scale range | [480, 800] |
| Mosaic probability | 0.5 |
| MixUp alpha | 0.2 |
| Copy-paste probability | 0.3 |

### v2

@@ VII Conclusion
This paper has presented MCAQ-YOLO, a morphological complexity-aware quantization framework that dynamically allocates bit-widths based on local visual complexity. The work makes two principal contributions beyond incremental accuracy improvements:
First, morphological complexity is established as a principled predictor of spatial quantization sensitivity ( at tile level, ), representing a shift from network-centric metrics (Hessian, activation statistics) to signal-centric metrics that directly characterize input difficulty. This perspective opens new directions for content-adaptive compression in visual recognition.
Second, it is demonstrated that tile-wise spatial mixed-precision quantization—long considered impractical for real-time inference—can be deployed efficiently through a calibration-time analysis design. On the construction safety dataset, MCAQ-YOLO achieves 85.6% mAP@0.5 with an average of 4.2 bits, outperforming uniform 4-bit quantization by 3.5 percentage points while maintaining 151 FPS. The modest gains on COCO (+2.9%) and VOC (+2.3%) indicate that spatial mixed-precision is most beneficial for high-variability deployment scenarios rather than curated benchmarks.
Future work. Future research directions include pursuing formal information-theoretic justification for the complexity–sensitivity relationship, extending the framework to Transformer-based detectors (DETR, RT-DETR), and developing hardware-native implementations for edge deployment.
Appendix A CUDA Kernel Implementation
The tile-wise mixed-precision quantization is realized through a custom CUDA kernel that performs per-element quantization with spatially varying bit-widths.
Kernel optimizations: (1) Memory coalescing for aligned access. (2) Bit map caching in shared memory. (3) Fused quantize-dequantize. (4) Vectorized loads (float4) for bandwidth.
Appendix B Extended Ablation Studies
Table VII presents comprehensive ablation results with extended metrics beyond mAP@0.5.
| Configuration | mAP@0.5 | mAP@[.5:.95] | APS | APM | APL | -value |
| Full MCAQ-YOLO | 85.60.4 | 63.10.4 | 36.30.5 | 67.00.4 | 80.70.4 | – |
| w/o Curriculum learning | 83.10.6 | 60.20.5 | 33.10.7 | 64.50.5 | 78.20.5 | |
| w/o Temperature annealing | 84.30.5 | 61.80.4 | 34.70.6 | 65.80.4 | 79.40.4 | |
| w/o Smoothness regularization | 84.20.5 | 61.50.4 | 34.20.6 | 65.30.4 | 79.10.4 | |
| Linear mapping (no MLP) | 83.80.4 | 60.90.4 | 33.80.5 | 64.90.4 | 78.80.4 | |
| w/o Fractal dimension () | 85.10.4 | 62.50.4 | 35.60.5 | 66.40.4 | 80.10.4 | |
| w/o Texture entropy () | 84.50.4 | 61.90.4 | 34.90.5 | 65.90.4 | 79.50.4 | |
| w/o Gradient variance () | 84.80.4 | 62.20.4 | 35.20.5 | 66.20.4 | 79.80.4 | |
| w/o Edge density () | 84.60.4 | 62.00.4 | 35.00.5 | 66.00.4 | 79.60.4 | |
| w/o Contour complexity () | 85.20.4 | 62.60.4 | 35.70.5 | 66.50.4 | 80.20.4 | |
| w/o Texture + Edge | 84.10.4 | 61.40.4 | 34.00.5 | 65.10.4 | 78.90.4 | |
| Single metric (Entropy only) | 83.40.5 | 60.50.5 | 33.30.6 | 64.30.5 | 78.40.5 | |
| w/o Knowledge distillation | 84.40.4 | 61.70.4 | 34.60.5 | 65.70.4 | 79.30.4 |
Tile Size Sensitivity. The default grid ( tiles) provides the best trade-off: 85.6% mAP with 1.8 ms overhead. Finer grids (: 85.9% mAP, 4.2 ms; : 86.0% mAP, 12.1 ms) yield marginal accuracy improvements at substantially higher cost.
Table VIII provides additional ablation results for tile size, complexity threshold, and bit range.
| Configuration | mAP@0.5 | FPS | |
| Tile Size Variations | |||
| Tile size = 16 | 85.4 | 0.2 | 143 |
| Tile size = 32 (default) | 85.2 | – | 151 |
| Tile size = 64 | 84.6 | 0.6 | 158 |
| Complexity Threshold | |||
| = 0.4 | 84.5 | 0.7 | 150 |
| = 0.6 (default) | 85.2 | – | 151 |
| = 0.8 | 84.7 | 0.5 | 152 |
| Bit Range | |||
| 3-5 bits | 84.1 | 1.1 | 156 |
| 3-6 bits (default) | 85.2 | – | 151 |
| 4-6 bits | 85.0 | 0.2 | 148 |
| Linear mapping only | 83.8 | 1.4 | 153 |
| Layer-wise only (no spatial) | 83.7 | 1.5 | 155 |
Appendix C Statistical Analysis Methodology
Bootstrap CIs. Confidence intervals are computed using the bias-corrected and accelerated (BCa) bootstrap method with 10,000 resamples.
Multiple comparisons. For ablation studies involving comparisons, we apply the Benjamini–Hochberg procedure to control the false discovery rate (FDR) at .
Power analysis. For the main comparison (MCAQ vs. Uniform 4-bit), with observed mean difference percentage points and pooled standard deviation (computed from the 10 independent runs), the effect size is . With runs per condition, the achieved power is approximately 95% at .
Appendix D Dataset Details
D-A Data Sources and Licensing
The CSE dataset is compiled from three publicly available PPE detection datasets, all under permissive licenses (CC BY 4.0 or Apache 2.0):
Safety Helmet Detection Dataset (Roboflow, CC BY 4.0): 4,500 images
PPE Detection Dataset (Kaggle, Apache 2.0): 5,200 images
Construction Site Safety Dataset (Open Images, CC BY 4.0): 2,800 images
Total raw images: 12,500. After cleaning (see below), 12,000 images remain.
D-B Cleaning and Deduplication Protocol
1.
Invalid annotation removal: Boxes with zero area, negative coordinates, or aspect ratios 20:1 were removed (135 boxes removed; images retained).
2.
Near-duplicate detection: Perceptual hashing (pHash) with Hamming distance 8 identified duplicate/near-duplicate images; 500 images removed to yield 12,000 unique images.
3.
Cross-source deduplication: Images appearing in multiple source datasets were counted only once.
D-C Train/Val/Test Split Protocol
The final 12,000 images are split into train (10,000), validation (1,000), and test (1,000):
Images were first grouped by source video/scene (where metadata was available).
Groups were randomly assigned to train (83.3%), val (8.3%), or test (8.3%).
No image from the same scene/video appears in multiple splits.
Final verification: pHash similarity between any train–test image pair exceeds Hamming distance 12.
D-D Class Statistics
Table IX provides per-class statistics for the final 12,000-image dataset.
| Class | Train | Val | Test | Total |
| Person | 12,450 | 1,245 | 1,248 | 14,943 |
| Helmet | 10,832 | 1,083 | 1,085 | 13,000 |
| Vest | 9,521 | 952 | 954 | 11,427 |
| Gloves | 6,234 | 623 | 625 | 7,482 |
| Boots | 5,892 | 589 | 591 | 7,072 |
| Goggles | 3,568 | 357 | 358 | 4,283 |
| Total | 48,497 | 4,849 | 4,861 | 58,207 |
Appendix E Morphological Metric Details
Fractal Dimension (). Estimated using multi-resolution box counting (Algorithm 2). The edge map is computed using Canny edge detection with adaptive thresholds (). The fractal dimension captures the self-similarity and geometric complexity of edge patterns, with higher values indicating more irregular boundaries.
Texture Entropy (). Computed from the histogram of uniform LBP codes:
where is the normalized frequency of LBP code , neighbors, and . Normalized by . Higher entropy indicates more diverse local texture patterns.
Gradient Variance (). Computed from Sobel gradients:
where and are horizontal and vertical gradient magnitudes computed via Sobel operators. This metric captures local contrast and edge strength.
Edge Density (). Ratio of edge pixels to total pixels in the tile:
where is the binary edge map from Canny detection with adaptive thresholds based on the Otsu method.
Contour Complexity (). Circularity-based measure averaged over detected contours:
where and are the perimeter and area of contour , and is the number of contours. A circle has circularity 1; more complex shapes have higher values.
Appendix F Detailed Hyperparameters
Table X provides complete hyperparameter specifications for reproducibility.
| Category | Parameter | Value |
| Optimizer | Type | AdamW |
| Learning rate | ||
| Weight decay | 0.05 | |
| 0.9 | ||
| 0.999 | ||
| Gradient clipping | 1.0 | |
| Learning Rate | Schedule | Cosine annealing |
| Warmup epochs | 5 | |
| Warmup LR | ||
| Minimum LR | ||
| Training | Total epochs | 300 |
| Batch size | 64 (16 4 GPUs) | |
| Input resolution | ||
| Random seeds | 10 runs | |
| Curriculum | Initial threshold | 0.2 |
| Warmup epochs | 20 | |
| Transition epochs | 20–50 | |
| Initial temperature | 10.0 | |
| Final temperature | 1.0 | |
| Loss Weights | (bit budget) | 0.01 0.1 |
| (smoothness) | 0.1 | |
| (KD) | 0.5 | |
| Bit target | 4.0 | |
| Quantization | Bit range | [2, 8] |
| Calibration images | 1,000 | |
| EMA momentum | 0.99 | |
| Default tile grid | ||
| Mapping MLP | Hidden dimensions | [32, 64, 32] |
| Activation | ReLU | |
| Normalization | BatchNorm | |
| Morphology | LBP radius | 1 |
| LBP neighbors | 8 | |
| Cache size | 10,000 entries | |
| Data Aug. | Mosaic probability | 0.5 |
| MixUp probability | 0.1 | |
| HSV augmentation | (0.015, 0.7, 0.4) | |
| Random flip | 0.5 |
Appendix G Energy Measurement Protocol
Energy consumption is measured using the NVIDIA Management Library (NVML) on an RTX PRO 6000 Blackwell Workstation Edition GPU with fixed clock frequencies (2617 MHz boost, 1750 MHz memory). Each session consists of 100 warm-up iterations followed by 500 timed iterations at batch size 1, with power sampled at 10 ms intervals and idle baseline subtracted. The reported energy (J) is computed as , averaged over 5 runs.
Appendix H Runtime Analysis Details
Table XI provides a more detailed runtime breakdown including optimization impact.
| Component | CPU | GPU | Opt. | Peak | Avg |
| (ms) | (ms) | (ms) | (MB) | (MB) | |
| YOLOv8 Backbone | 15.2 | 4.8 | 4.8 | 45.2 | 32.1 |
| Morphological Analysis | 12.8 | 3.5 | 1.4 | 31.7 | 22.0 |
| Bit Allocation + Quant. | 3.6 | 1.0 | 0.4 | 3.3 | 2.3 |
| Total Overhead | 16.4 | 4.5 | 1.8 | 35.0 | 24.3 |
| Total with YOLO | 31.6 | 9.3 | 6.6 | 76.9 | 54.1 |
Appendix I Limitations and Future Work
1.
Empirical foundation: The complexity–sensitivity relationship is empirically established; formal information-theoretic justification remains open.
2.
Hardware deployment: Custom CUDA kernel required; native hardware support for spatially varying precision is lacking.
3.
Dataset dependency: Performance gains correlate with dataset morphological variability.
4.
Architecture scope: Validation limited to CNN-based YOLO; Transformer-based architectures require adaptation.
Future directions: Theoretical analysis of complexity–sensitivity relationships, Transformer extension, hardware co-design, and joint compression with pruning/NAS.